A sodium atom has a mass of 22.99 u. Use Avogadro’s

constant and the conversion factor 1.661 3 10–24 g/u to

calculate the mass of 1.00 mol of sodium atoms.


1
Expert's answer
2022-06-15T10:57:04-0400

1 atom Na = 22.99 u

1 u = 1.661*10-24 g

1atom Na=22.99* 1.661*10-24 =3.82*10-23 g

According to Avogadro's law, 1 mole of a substance contains 6.02 * 1023 particles (atoms, molecules).

"1~mol ~Na = ~6.02*~10^{23} * ~3.82*~10^{-23} =~\bold{23 ~g}"


Answer: 23 g